### v1

# latexText page

Abstract
Model-based iterative reconstruction plays a key role in solving inverse problems. However, the associated minimization problems are generally large-scale, ill-posed, nonsmooth, and sometimes even nonconvex, which present challenges in designing efficient iterative solvers and often prevent their practical use. Preconditioning methods can significantly accelerate the convergence of iterative methods. In some applications, computing preconditioners on-the-fly is beneficial. Moreover, forward models in image reconstruction are typically represented as operators, and the corresponding explicit matrices are often unavailable, which brings additional challenges in designing preconditioners. Therefore, for practical use, computing and applying preconditioners should be computationally inexpensive. This paper adapts the randomized Nyström approximation to compute effective preconditioners that accelerate image reconstruction without requiring an explicit matrix for the forward model. We leverage modern GPU computational platforms to compute the preconditioner on-the-fly. Moreover, we propose efficient approaches for applying the preconditioner to problems with nonsmooth regularizers. Our numerical results on image deblurring, super-resolution with impulsive noise, and computed tomography reconstruction demonstrate the efficiency and effectiveness of the proposed preconditioner.
Index Terms:
image deblur, super-resolution, CT reconstruction, Nyström preconditioner, wavelet, total variation, Hessian Schatten-norm.
I Introduction
THE task of image reconstruction is to recover a clean image from degraded measurements. Model-based iterative reconstruction recovers a clean image $\bm{\mathrm{x}}$ by solving the following type of minimization problem:
$\min_{\bm{\mathrm{x}}\in\mathbb{R}^{N}}f(\bm{\mathrm{x}})+\lambda\,g(\bm{\mathrm{x}}),$ (1)
where $f(\bm{\mathrm{x}})$ denotes a data fidelity specifying the discrepancy between the model’s predictions and the degraded measurements. $g(\bm{\mathrm{x}})$ denotes a regularizer imposing prior information on $\bm{\mathrm{x}}$. $\lambda>0$ is a trade-off parameter to balance $f(\bm{\mathrm{x}})$ and $g(\bm{\mathrm{x}})$. Both $f(\bm{\mathrm{x}})$ and $g(\bm{\mathrm{x}})$ can be nonsmooth and nonconvex. This paper focuses on linear inverse problems where we solve
$\min_{\bm{\mathrm{x}}\in\mathbb{R}^{N}}\frac{1}{p}\underbrace{\|\bm{\mathrm{A}}\bm{\mathrm{x}}-\bm{\mathrm{y}}\|_{p}^{p}}_{f(\bm{\mathrm{x}})}+\lambda\,g(\bm{\mathrm{x}}),$ (2)
where $\bm{\mathrm{A}}\in\mathbb{R}^{M\times N}$ with $M\leq N$ is called the forward model describing the mapping from the image of interest $\bm{\mathrm{x}}$ to the measurements $\bm{\mathrm{y}}$. In this paper, we assume $p\in(0,2]$.
In practice, the matrix $\bm{\mathrm{A}}$ is under-determined or ill-conditioned and the measurements are noisy, so one uses a regularizer $g(\bm{\mathrm{x}})$ to stabilize the solution. Typical choices of $g(\bm{\mathrm{x}})$ that are effective include wavelet [1], total variation (TV) [2], dictionary learning [3], low-rank [4]. When using such hand-crafted regularizers, we refer to (1) as variational image reconstruction. Over the past decades, deep learning (DL) [5] has received significant attention in the field of image reconstruction. Instead of using hand-crafted priors, DL-based methods learn implicit priors from massive training data. This includes end-to-end learning [6] and physics-informed deep unrolling [7, 8]. Alternatively, one can use plug-and-play [9, 10, 11] or regularization by denoising [12, 13] where one trains a deep denoiser [14] first and then uses it as an implicit prior for general inverse problems. Recently, building priors with generative models, such as diffusion models, has also attracted much attention. We refer the reader to [15] for a survey of those directions.
While deep learning methods often outperform variational methods in many applications, variational image reconstruction [16, 17, 18, 19] offers significant advantages in terms of interpretability, theoretical guarantees, and robustness to noise or distribution shifts. Variational methods do not require large training datasets and provide predictable, stable behavior. Moreover, they allow for easier customization of priors based on domain knowledge, making them particularly suitable for applications where interpretability and reliability are critical, such as medical imaging. Recent work has also shown that novel neural network structures can be derived from variational methods, yielding state-of-the-art performance [20, 21, 22, 23, 24]. Therefore, this paper considers hand-crafted regularizers for (2). Specifically, we consider $g(\bm{\mathrm{x}})=\Psi(\bm{\mathrm{L}}\bm{\mathrm{x}})$ where $\bm{\mathrm{L}}\in\mathbb{R}^{L\times N}$ refers to the regularization operator and $\Psi(\cdot)$ denotes an energy function. A typical choice of $\Psi(\cdot)$ is a norm or quasi-norm.
I-A $\ell_{p}-\ell_{q}$ Image Reconstruction
For $p\in(0,2]$, we consider $\Psi(\cdot)=\frac{1}{q}\|\cdot\|_{q}^{q}$ with $q\in(0,2]$ and (2) becomes
$\min_{\bm{\mathrm{x}}\in\mathbb{R}^{N}}\underbrace{\frac{1}{p}\|\bm{\mathrm{A}}\bm{\mathrm{x}}-\bm{\mathrm{y}}\|_{p}^{p}}_{f(\bm{\mathrm{x}})}+\lambda\underbrace{\frac{1}{q}\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{q}^{q}}_{g(\bm{\mathrm{x}})},$ (3)
which is known as the $\ell_{p}-\ell_{q}$ problem [16]. For $p,q\leq 1$, both $f(\bm{\mathrm{x}})$ and $g(\bm{\mathrm{x}})$ are nonsmooth. To address the nonsmooth challenge, Clason et al. considered (3) in its dual formulation [25], where the objective function becomes differentiable, at the cost of additional box and linear equality constraints. For general $p,\,q\in(0,2)$, Huang et al. [26] first smoothed $f$ and $g$ and then applied the majorization-minimization for the smoothed problem, leading to an inner least-squares problem. Then [26] used the Krylov based methods to solve the smoothed problem for acceleration. Following the same strategy, Buccini et al. [27] applied modulus-based iterative methods [28] to the smoothed problem to address box constraints. Alternatively, Chan et al. [29] solved (3) with a half-quadratic algorithm. Lanza et al. [30] proposed a Krylov subspace method to reduce computation of the method in [29] by assuming that the signal can be represented in a subspace.
Such solvers for (3) can be considered as iteratively reweighted methods (IRM) [16] because we eventually need to solve a linear equation (19) with updating of the associated weightings at each iteration. For completeness, Section III-A describes one type IRM for (3). Typically the appealing conjugate gradient (CG) method is chosen to address the associated linear equation. However, the convergence of CG can be extremely slow for some problems, especially for $p,\,q<1$, as the problem becomes nonconvex. Preconditioning methods [31, 32] are widely used to accelerate the convergence of CG. Because the weightings $\bm{\mathrm{W}}_{f}^{k}$ and $\bm{\mathrm{W}}_{g}^{f}$ (defined in (17)) change at each iteration, an effective preconditioner should change at each iteration, so computing such a preconditioner must be computationally inexpensive.
I-B $\ell_{2}$ with Mixed Norm Image Reconstruction
When $p=2$, we consider $\Psi(\cdot)$ to be an $\ell_{1,\phi}$ mixed norm (i.e., $\|\mathcal{V}\|_{1,\phi}=\sum_{l=1}^{G}\|\bm{\mathrm{v}}_{l}\|_{\phi}$ with $\phi\geq 1$ and $\bm{\mathrm{v}}_{l}$ denoting the $l$th group of $\mathcal{V}=\{\bm{\mathrm{v}}_{1},\bm{\mathrm{v}}_{2},\cdots,\bm{\mathrm{v}}_{G}\}$), for which we can develop more efficient approaches than IRM to solve (3) and a broader range of choices for $\bm{\mathrm{L}}$. Moreover, we can also address convex constraints $\mathcal{C}$ easily. Dividing the components of $\mathcal{V}$ into $G$ groups, the mixed norm $\ell_{\psi,\phi}$ is defined by applying an $\ell_{\phi}$ norm to each group and then applying an $\ell_{\psi}$ norm to the entire set of groups. In this paper, we assume $\psi=1$. Specifically, we address
$\min_{\bm{\mathrm{x}}\in\mathcal{C}}\underbrace{\frac{1}{2}\|\bm{\mathrm{A}}\bm{\mathrm{x}}-\bm{\mathrm{y}}\|_{2}^{2}}_{f(\bm{\mathrm{x}})}+\lambda\underbrace{\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{1,\phi}}_{g(\bm{\mathrm{x}})}.$ (4)
When $\bm{\mathrm{L}}$ is a first-order differential operator and $\phi=1$ (respectively, $\phi=2$), then (4) represents the anisotropic (respectively, isotropic) TV-based reconstruction. When $\bm{\mathrm{L}}$ denotes the second-order finite-difference operator, then (4) becomes the Hessian Schatten (HS) norm based reconstruction [33]. Appendix A presents an explicit form for $\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{1,\phi}$ for various $\bm{\mathrm{L}}$ to clarify the definition of mixed norm.
Because (4) is a standard composite minimization problem, an appealing algorithm for solving it is the accelerated proximal gradient (APG) method [34]. Alternatively, one could use a primal-dual method [35, 36] or ADMM [37] etc. To accelerate the convergence of solving (4), one can adapt a quasi-Newton proximal method that uses second-order information [38, 39, 40]. However, the efficiency of using a quasi-Newton proximal method depends highly on the accuracy of approximating the Hessian matrix of $f(\bm{\mathrm{x}})$ and the efficiency of solving the associated weighted proximal mapping (WPM, defined in (5)).
I-B1 Previous Work on Preconditioning Methods in Image Reconstruction
Preconditioning methods are used widely to accelerate the convergence of iterative methods [41] in scientific computing and have been extended to accelerate the convergence of image reconstruction. Clinthorne et al. [42] applied a spatially invariant preconditioner to accelerate the convergence of reconstructing single-photon emission computed tomography (SPECT) images with quadratic $g(\bm{\mathrm{x}})$. Fessler et al. [43] proposed a shift-variant preconditioner for accelerating positron emission tomography (PET) reconstruction with non-quadratic smooth $g(\bm{\mathrm{x}})$. Lin et al. [44] derived cosine transform and incomplete factorization based preconditioners for image super-resolution task with smooth $g(\bm{\mathrm{x}})$. Pelletier et al. [45] developed a block circulant with circulant block (BCCB) preconditioner for edge-preserving image super-resolution. The authors in [46, 47] considered diagonal preconditioners for accelerating PET and magnetic resonance imaging reconstruction. To reduce the computation of computing WPM, [48] discussed an unmatched preconditioner method for image reconstruction by computing WPM approximately. To address the nonsmoothness of $g(\bm{\mathrm{x}})$, Koolstra et al. [49] adapted the preconditioner to ADMM to accelerate the convergence of solving the associated least-squares problem. By leveraging deep learning, [50, 51] proposed methods to learn preconditioners to accelerate the convergence of image reconstruction.
I-C Our Contribution
The preconditioners discussed in Section I-B1 were either used for smooth $g(\bm{\mathrm{x}})$ [42, 43, 44, 46, 47] or applied to a nonsmooth $g(\bm{\mathrm{x}})$ after smoothing [45, 51]. Moreover, the previous methods require domain knowledge because they were developed for specific applications, making them difficult to generalize to other applications. Although the preconditioners proposed in [49, 50] can be used to address nonsmooth $g(\bm{\mathrm{x}})$, the convergence rate of ADMM is only $\mathcal{O}(1/k)$. To address these challenges, we adapt the randomized Nyström approximation (RNA) to design an effective preconditioner, called randomized Nyström preconditioner (RNP), to accelerate variational image reconstruction. Our main contributions are summarized as follows.
[…]
  • By leveraging modern high-performance GPU computational platforms, we compute RNP on-the-fly, allowing it to accelerate IRM for (3) in a computationally efficient way. Our numerical experiments on image deblurring and super-resolution with impulsive noise showed that, by using RNP, we can reduce wall time by more than $90$%. To the best of our knowledge, this is the first work to consider the use of hardware platform in conjunction with computing preconditioners on-the-fly in image reconstruction.
  • We show how RNP can be adapted to directly address nonsmooth $g(\bm{\mathrm{x}})$. Specifically, we propose efficient approaches for integrating RNP with APG to solve (4). Our experiments on computed tomography (CT) reconstruction with wavelet, TV, and HS norm regularizers demonstrate that RNP can significantly accelerate the convergence of APG.
I-D Preliminaries and Roadmap
This subsection introduces some definitions and a theorem that will be used frequently in the subsequent discussion, and then presents a roadmap of this paper.
Definition 1 (Weighted proximal mapping (WPM)).
Given a proper closed convex function $h(\bm{\mathrm{x}})$, a symmetric positive-definite matrix $\bm{\mathrm{W}}\succ 0\in\mathbb{R}^{N\times N}$, and $\lambda>0$, the WPM associated with $h$ is defined as
$\mathrm{prox}_{\lambda h}^{\bm{\mathrm{W}}}(\bm{\mathrm{x}})=\arg\min\limits_{\bm{\mathrm{u}}\in\mathbb{R}^{N}}\left(\lambda h(\bm{\mathrm{u}})+\frac{1}{2}\|\bm{\mathrm{u}}-\bm{\mathrm{x}}\|^{2}_{\bm{\mathrm{W}}}\right),$ (5)
where $\|\bm{\mathrm{x}}\|_{\bm{\mathrm{W}}}\triangleq{\sqrt{\bm{\mathrm{x}}^{\mathrm{T}}\bm{\mathrm{W}}\bm{\mathrm{x}}}}$ denotes the $\bm{\mathrm{W}}$-norm.
For $\bm{\mathrm{W}}=\bm{\mathrm{I}}$, (5) simplifies to the proximal mapping:
$\mathrm{prox}_{\lambda h}(\bm{\mathrm{x}})=\arg\min_{\bm{\mathrm{u}}\in\mathbb{R}^{N}}\lambda h(\bm{\mathrm{u}})+\frac{1}{2}\|\bm{\mathrm{u}}-\bm{\mathrm{x}}\|_{2}^{2}.$ (6)
If $h(\bm{\mathrm{x}})$ is a characteristic function
$\delta_{\mathcal{C}}(\bm{\mathrm{x}})=\left\{\begin{array}[]{rl}0&\bm{\mathrm{x}}\in\mathcal{C}\\ \infty&\bm{\mathrm{x}}\notin\mathcal{C},\end{array}\right.$
then the proximal mapping becomes projection onto the set $\mathcal{C}$.
In general, computing a WPM (5) is computationally expensive. However, when $\bm{\mathrm{W}}$ is represented by a diagonal matrix plus a rank-$r$ correction, then we can use Theorem 1 to compute WPM relatively cheaply.
Theorem 1 ([39], Theorem 3.4).
Let $\bm{\mathrm{W}}=\bm{\mathrm{D}}\pm{\bm{\mathrm{U}}}{\bm{\mathrm{U}}}^{\mathsf{T}}$, $\bm{\mathrm{W}}\succ 0\in\mathbb{R}^{N\times N}$, and $\bm{\mathrm{U}}\in\mathbb{R}^{N\times r}$. Then, it holds that
$\mathrm{prox}_{\lambda h}^{\bm{\mathrm{W}}}(\bm{\mathrm{x}})=\mathrm{prox}_{\lambda h}^{\bm{\mathrm{D}}}(\bm{\mathrm{x}}\mp{\bm{\mathrm{D}}}^{-1}{\bm{\mathrm{U}}}\bm{\gamma}^{*}),$ (7)
where $\bm{\gamma}^{*}\in\mathbb{R}^{r}$ is the unique solution of
${\bm{\mathrm{U}}}^{\mathsf{T}}\left(\bm{\mathrm{x}}-\mathrm{prox}_{\lambda h}^{\bm{\mathrm{D}}}\left(\bm{\mathrm{x}}\mp{\bm{\mathrm{D}}}^{-1}{\bm{\mathrm{U}}}\bm{\gamma}\right)\right)+\bm{\gamma}=\bm{0}.$ (8)
Semi-smooth Newton methods can solve (8) efficiently [52, 39, 53].
The rest of this paper is organized as follows. Section II describes how RNA can be a preconditioner for linear equations. Section III shows how to adapt RNP to accelerate IRM and APG for solving (3) and (4), respectively. Section IV presents numerical experiments on image deblurring and super-resolution with impulsive noise and computed tomography reconstruction to demonstrate the effectiveness and efficiency of using RNP for acceleration. Section V presets the conclusion and future work.
II Randomized Nyström Approximation
The Nyström method provides a way to build a low-rank approximation for a symmetric positive semidefinite (SPSD) matrix $\bm{\Phi}\succeq 0\in\mathbb{R}^{N\times N}$. Let $\bm{\Omega}\in\mathbb{R}^{N\times K}$ be a random test matrix [54, 55] with sketch size $K\geq 1$. The Nyström method uses the following formula as a low-rank approximation of $\bm{\Phi}$, called the Nyström approximation
$\bm{\Phi}\left<\bm{\Omega}\right>=(\bm{\Phi}\bm{\Omega})(\bm{\Omega}^{\mathrm{T}}\bm{\Phi}\bm{\Omega})^{\dagger}(\bm{\Phi}\bm{\Omega})^{\mathrm{T}},$ (9)
where denotes the pseudo-inverse. Clearly $\bm{\Phi}\left<\bm{\Omega}\right>$ is also a SPSD matrix and its rank is at most $K$ [56, Lemma A.1]. Because a direct implementation of (9) is numerically unstable, we use the robust implementation proposed in [54], summarized in Algorithm 1. That algorithm computes $\bm{\mathrm{U}}$ and $\hat{\bm{\mathrm{S}}}$ such that $\bm{\Phi}\left<\bm{\Omega}\right>=\bm{\mathrm{U}}\hat{\bm{\mathrm{S}}}\bm{\mathrm{U}}^{\mathrm{T}}$, although in practice we do not compute that product explicitly.
Many scientific computing applications [57] require solving the following linear equation:
$(\bm{\Phi}+\mu\bm{\mathrm{I}})\bm{\mathrm{x}}=\bm{\mathrm{b}},~{}\mu>0.$ (10)
Using RNA to formulate a preconditioner to address (10), the associated formulations of RNP and its inverse [55] are as follows:
$\begin{array}[]{ll}\bm{\mathrm{P}}\quad=&\frac{1}{{\hat{s}}_{K}+\mu}\bm{\mathrm{U}}(\hat{\bm{\mathrm{S}}}+\mu\bm{\mathrm{I}})\bm{\mathrm{U}}^{\mathrm{T}}+(\bm{\mathrm{I}}-\bm{\mathrm{U}}\bm{\mathrm{U}}^{\mathrm{T}}),\\ \bm{\mathrm{P}}^{-1}=&({\hat{s}}_{K}+\mu)\bm{\mathrm{U}}(\hat{\bm{\mathrm{S}}}+\mu\bm{\mathrm{I}})^{-1}\bm{\mathrm{U}}^{\mathrm{T}}+(\bm{\mathrm{I}}-\bm{\mathrm{U}}\bm{\mathrm{U}}^{\mathrm{T}}),\end{array}$ (11)
where ${\hat{s}}_{K}$ is the $K$th eigenvalue of $\bm{\Phi}\left<\bm{\Omega}\right>$. $\bm{\mathrm{U}}$ and $\hat{\bm{\mathrm{S}}}$ are defined in Algorithm 1. Let $d_{\mathrm{eff}}(\mu)=\mathrm{tr}\left(\bm{\Phi}(\bm{\Phi}+\mu\bm{\mathrm{I}})^{-1}\right)$ be the effective dimension. Then, we have the following theorem to describe the condition number of $(\bm{\Phi}+\mu\bm{\mathrm{I}})$ after using RNP, which reveals the effectiveness of RNP.
Theorem 2 ([55], Theorem 5.1).
Suppose we build the Nyström preconditioner through (11) with sketch size $K=2\lceil 1.5d_{\mathrm{eff}}(\mu)+1\rceil$. Then using $\bm{\mathrm{P}}$ as the preconditioner for $(\bm{\Phi}+\mu\bm{\mathrm{I}})$ results in the expectation of condition number bound
$\mathbb{E}[\kappa(\bm{\mathrm{P}}^{-1/2}(\bm{\Phi}+\mu\bm{\mathrm{I}})\bm{\mathrm{P}}^{-1/2})]<28,$
where $\kappa(\cdot)$ denotes the condition number function.
Theorem 2 shows that, by choosing an appropriate sketch size $K$, RNP can significantly reduce the condition number of $(\bm{\Phi}+\mu\bm{\mathrm{I}})$. Thus, one can significantly accelerate the convergence of iterative methods for solving (10) with RNP. However, in practice, computing $d_{\mathrm{eff}}$ is computationally expensive. Moreover, the value of $K=2\lceil 1.5d_{\mathrm{eff}}(\mu)+1\rceil$ can be very large, which will significantly increases the computation of computing and applying RNP. In our experimental settings, we found that using a moderate value for $K$ is sufficient. We refer the reader to [55, 54], where the authors discussed more theoretical properties of using RNP for considering (10). The main purpose of this paper is to study how to adapt RNP to accelerate variational image reconstruction, focusing on evaluating its effectiveness and efficiency. We also refer the reader to [55, 56, 58] and the references therein that discuss using RNP for other applications.
Algorithm 1 Nyström approximation.
0:  Sketch size $K$, machine accuracy $\epsilon$, and symmetric positive semidefinite operator $\bm{\Phi}\in\mathbb{R}^{N\times N}$.
0:  $\bm{\mathrm{U}}\in\mathbb{R}^{N\times K},~{}\hat{\bm{\mathrm{S}}}\in\mathbb{R}^{K\times K}$.
1:  Generate a random matrix $\bm{\Omega}\in\mathbb{R}^{N\times K}$
2:  $\bm{\mathrm{Y}}=\bm{\Phi}\bm{\Omega}\in\mathbb{R}^{N\times K}$
3:  $v=\epsilon\|\bm{\Omega}\|_{F}$
4:  $\bm{\mathrm{Y}}_{v}=\bm{\mathrm{Y}}+v\bm{\Omega}$
5:  $\bm{\mathrm{C}}=\mathrm{chol}({\mathbf{\Omega}}^{\mathrm{T}}\bm{\mathrm{Y}}_{v})\in\mathbb{R}^{K\times K}$ Cholesky decomp.
6:  $\bm{\mathrm{B}}=\bm{\mathrm{Y}}_{v}\bm{\mathrm{C}}^{-{\mathrm{T}}}\in\mathbb{R}^{N\times K}$
7:  $[\bm{\mathrm{U}},\bm{\mathrm{S}},\sim]=\mathrm{SVD}(\bm{\mathrm{B}})$
8:  $\hat{\bm{\mathrm{S}}}=\max\{\bm{0},\bm{\mathrm{S}}^{2}-v\bm{\mathrm{I}}\}$
III On Adapting Nyström Preconditioners in Image Reconstruction
This section first describes IRM in Section III-A and then shows how RNP can be adapted to accelerate IRM for solving (3) efficiently. Because the weightings in IRM change at each iteration, computing RNP must be computationally inexpensive. We leverage the modern GPU computational platform and the PyTorch library to present an on-the-fly implementation for computing RNP at the end of this section. Additionally, we show how RNP can be incorporated with APG [59] to efficiently handle (4). Furthermore, we propose efficient approaches to solve the associated WPM to reduce the computational cost involved in applying RNP. We consider separately the $\ell_{p}-\ell_{q}$ and $\ell_{2}-\ell_{1,\phi}$ cases.
III-A $\ell_{p}-\ell_{q}$ Image Reconstruction Acceleration
Following [29, Lemma 1], we represent the $p$th power of $r\in\mathbb{R}\setminus\{0\}$ with $20<p<$ 2 as
$|r|^{p}=\min_{\beta>0}\left\{\beta r^{2}+\frac{1}{b_{p}\beta^{a_{p}}}\right\},$ (12)
where $a_{p}=\frac{p}{2-p}$ and $b_{p}=\frac{2^{\frac{2}{2-p}}}{(2-p)p^{\frac{2}{2-p}}}$ are two positive scalars. The minimizer of (12) is
$\beta^{*}=\frac{p}{2}|r|^{p-2}.$ (13)
For the singular case, $r=0$, we simply set $|r|_{\epsilon}=\sqrt{r^{2}+\epsilon}$ with $\epsilon>0$, following [29].
Representing (3) in a component-wise form, we get
$\min_{\bm{\mathrm{x}}}\frac{1}{p}\sum_{m=1}^{M}|(\bm{\mathrm{A}}\bm{\mathrm{x}})_{m}-y_{m}|^{p}+\frac{\lambda}{q}\sum_{l=1}^{L}\sum_{\zeta=1}^{\pi}|(\bm{\mathrm{L}}\bm{\mathrm{x}})_{l}^{\zeta}|^{q},$ (14)
where $(\bm{\mathrm{A}}\bm{\mathrm{x}})_{m}$ and $y_{m}$ represent the $m$th element of $\bm{\mathrm{A}}\bm{\mathrm{x}}$, $\bm{\mathrm{y}}$, respectively. $(\bm{\mathrm{L}}\bm{\mathrm{x}})_{l}^{\zeta}$ denotes the $\zeta$th element in the $l$th group of $\bm{\mathrm{L}}\bm{\mathrm{x}}$. Here, we assume each group has the same number of elements and consider cases where each group represents a scalar or vector. Using (12), we rewrite (14) as
$\begin{array}[]{ll}\min\limits_{\begin{subarray}{c}\bm{\mathrm{x}},\\ \bm{\mathrm{v}}>\bm{0},\\ \bm{\mathrm{z}}>\bm{0}\end{subarray}}&F(\bm{\mathrm{x}},\bm{\mathrm{v}},\bm{\mathrm{z}})\equiv\frac{1}{p}\sum_{m=1}^{M}\Big{(}v_{m}|(\bm{\mathrm{A}}\bm{\mathrm{x}})_{m}-y_{m}|^{2}+\\[6.0pt] &\quad\quad\frac{1}{b_{p}v_{m}^{a_{p}}}\Big{)}+\frac{\lambda}{q}\sum_{l=1}^{L}\sum_{\zeta}\Big{(}z_{l}^{\zeta}|(\bm{\mathrm{L}}\bm{\mathrm{x}})_{l}^{\zeta}|^{2}+\frac{1}{b_{q}(z_{l}^{\zeta})^{a_{q}}}\Big{)},\end{array}$ (15)
where $v_{m}$ and $z_{l}^{\zeta}$ denote the $m$ and $(l-1)\pi+\zeta$th element of $\bm{\mathrm{v}}$ and $\bm{\mathrm{z}}$, respectively. By using an alternating minimization algorithm, IRM [16] solves (15) by successively updating $\bm{\mathrm{v}}$, $\bm{\mathrm{z}}$, and $\bm{\mathrm{x}}$. Algorithm 2 summarizes the IRM steps. Using (13), we derive closed-form solutions for $\bm{\mathrm{v}}^{k}$ and $\bm{\mathrm{z}}^{k}$ at steps 2 and 3 of Algorithm 2, i.e.,
$\begin{array}[]{rcl}\bm{\mathrm{v}}^{k}&=&\frac{p}{2}|(\bm{\mathrm{A}}\bm{\mathrm{x}}^{k})-\bm{\mathrm{y}}|^{p-2},\\[5.0pt] \bm{\mathrm{z}}^{k}&=&\frac{q}{2}|\bm{\mathrm{L}}\bm{\mathrm{x}}^{k}|^{q-2},\end{array}$ (16)
where the powers $(p-2)$ and $(q-2)$ are element-wise.
By defining
$\bm{\mathrm{W}}_{f}^{k}=\mathrm{diag}\left(\frac{2}{p}\bm{\mathrm{v}}^{k}\right)~{}~{}\text{and}~{}~{}\bm{\mathrm{W}}_{g}^{k}=\mathrm{diag}\left(\frac{2}{q}\bm{\mathrm{z}}^{k}\right),$ (17)
the minimization problem at step 4 of Algorithm 2 becomes
$\bm{\mathrm{x}}^{k}=\arg\min_{\bm{\mathrm{x}}\in\mathbb{R}^{N}}\|\bm{\mathrm{A}}\bm{\mathrm{x}}-\bm{\mathrm{y}}\|^{2}_{\bm{\mathrm{W}}_{f}^{k}}+\lambda\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{\bm{\mathrm{W}}_{g}^{k}}^{2}.$ (18)
The first-order optimality condition of (18) is given by
$\bm{\Phi}_{k}\bm{\mathrm{x}}=\bm{\mathrm{A}}^{\mathrm{T}}\bm{\mathrm{W}}_{f}^{k}\bm{\mathrm{y}},\quad\bm{\Phi}_{k}=\bm{\mathrm{A}}^{\mathrm{T}}\bm{\mathrm{W}}_{f}^{k}\bm{\mathrm{A}}+\lambda\bm{\mathrm{L}}^{\mathrm{T}}\bm{\mathrm{W}}_{g}^{k}\bm{\mathrm{L}}.$ (19)
Since the Hessian $\bm{\Phi}_{k}$ is a SPSD matrix, we use CG to solve (19). However, solving (19) is time-consuming when $\bm{\Phi}_{k}$ is ill-conditioned. we use (11) to construct a preconditioner for $\bm{\Phi}_{k}$ and apply the preconditioned CG (PCG) method to solve (19). Because $\bm{\mathrm{W}}_{f}^{k}$ and $\bm{\mathrm{W}}_{g}^{k}$ change at each iteration, computing $\bm{\mathrm{P}}$ in (11) must be computationally inexpensive. Section III-C discusses an on-the-fly implementation of RNP to address this computational issue. If $p,q=2$, then (3) is a quadratic minimization problem, allowing us to directly apply PCG with RNP. When $p=2,q\neq 1$, IRM is still applicable, but APG is more efficient as the following subsection discusses.
Algorithm 2 Alternating minimization method for solving (15).
0:  Initialization $\bm{\mathrm{x}}^{0}$, tolerance $\varepsilon$, and maximum number of iterations Max_Iter.
0:  $\bm{\mathrm{x}}^{*}$.
1:  for $k=1,2,\cdots,$ Max_Iter do
2:     $\bm{\mathrm{v}}^{k}=\arg\min_{\bm{\mathrm{v}}>0}F(\bm{\mathrm{x}}^{k-1},\bm{\mathrm{v}})$
3:     $\bm{\mathrm{z}}^{k}=\arg\min_{\bm{\mathrm{z}}>0}F(\bm{\mathrm{x}}^{k-1},\bm{\mathrm{z}})$
4:     $\bm{\mathrm{x}}^{k}=\arg\min_{\bm{\mathrm{x}}}F(\bm{\mathrm{x}},\bm{\mathrm{v}}^{k},\bm{\mathrm{z}}^{k})$
5:     if $\frac{\|\bm{\mathrm{x}}^{k}-\bm{\mathrm{x}}^{k-1}\|}{\|\bm{\mathrm{x}}^{k}\|}<\varepsilon$ then
6:        break
7:     end if
8:  end for
9:  $\bm{\mathrm{x}}^{*}=\bm{\mathrm{x}}^{k}$
III-B $\ell_{2}-\ell_{1,\phi}$ Image Reconstruction Acceleration
APG, which only uses first-order information and achieves the optimal worst-case convergence rate of $\mathcal{O}(1/k^{2})$ [60], is a very appealing algorithm for addressing (4). To further accelerate the convergence of APG, we adapt RNP to APG, resulting in the weighted accelerated proximal gradient (WAPG) algorithm that is summarized in Algorithm 3. For any $\bm{\mathrm{x}}_{1},\bm{\mathrm{x}}_{2}\in\mathbb{R}^{N}$, the Lipschitz constant $L_{f}^{\bm{\mathrm{P}}}$ of $f$ with respect to the $\bm{\mathrm{P}}$-norm satisfies the following inequality
$\|\bm{\mathrm{P}}^{-1}(\nabla f(\bm{\mathrm{x}}_{1})-\nabla f(\bm{\mathrm{x}}_{2}))\|_{\bm{\mathrm{P}}}\leq L_{f}^{\bm{\mathrm{P}}}\|\bm{\mathrm{x}}_{1}-\bm{\mathrm{x}}_{2}\|_{\bm{\mathrm{P}}}.$ (20)
If $f,g$ are convex, the worst-case convergence rate of the cost function sequence generated by Algorithm 3 is $\frac{2L_{f}^{\bm{\mathrm{P}}}\|\bm{\mathrm{x}}^{1}-\bm{\mathrm{x}}^{*}\|_{\bm{\mathrm{P}}}^{2}}{(k+1)^{2}}$ [61, 10.7.5] where $\bm{\mathrm{x}}^{1}$ and $\bm{\mathrm{x}}^{*}$ are the initial and global values, respectively. Clearly, for $\bm{\mathrm{P}}=\bm{\mathrm{I}}$, WAPG simplifies to APG. By choosing a suitable $\bm{\mathrm{P}}$, one can expect a lower bound than $\bm{\mathrm{P}}=\bm{\mathrm{I}}$ yielding faster convergence. Indeed, our experimental results in CT reconstruction illustrate the faster convergence of WAPG with RNP.
For noninvertible $\bm{\mathrm{L}}$, the associated WPM in Algorithm 3 generally does not have a closed-form solution, even with $\bm{\mathrm{P}}=\bm{\mathrm{I}}$, so WPM requires iterative methods. Following [53], we solve WPM using its dual problem to address the nonsmoothness of $g(\bm{\mathrm{x}})$. Moreover, compared with computing WPM with $\bm{\mathrm{P}}=\bm{\mathrm{I}}$, we show that solving WPM with RNP negligibly increases computation. Whereas [53] only discussed constrained TV regularization, we generalize their derivation to an abstract $\bm{\mathrm{L}}$ with mixed norm in this work.
Computing WPM at step 2 of Algorithm 3 at $k$th iteration is equivalent to solving
$\min_{\bm{\mathrm{x}}\in\mathcal{C}}\frac{1}{2}\|\bm{\mathrm{x}}-\bm{\mathrm{s}}^{k}\|_{\bm{\mathrm{P}}}^{2}+\bar{\lambda}\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{1,\phi},$ (21)
where $\bm{\mathrm{s}}^{k}=\bm{\mathrm{u}}^{k-1}-\alpha\bm{\mathrm{P}}^{-1}\nabla f(\bm{\mathrm{u}}^{k-1})$, $\bar{\lambda}=\alpha\lambda$, and $\alpha$ represents the stepsize. By using the fact that
$\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{1,\phi}=\max_{\mathcal{Q}\in B_{\infty,\psi}}\left\langle\mathcal{Q},\bm{\mathrm{L}}\bm{\mathrm{x}}\right\rangle,$ (22)
where $B_{\infty,\psi}$ represents the $\ell_{\infty,q}$ mixed unit-norm ball, i.e.,
$\begin{array}[]{rl}B_{\infty,\psi}=&\Big{\{}\mathcal{Q}=\{\bm{\mathrm{q}}_{1},\bm{\mathrm{q}}_{2},\ldots,\bm{\mathrm{q}}_{G}\}:\\ &\quad\|\bm{\mathrm{q}}_{l}\|_{\psi}\leq 1,~{}\forall l=1,\ldots,G\Big{\}},\end{array}$ (23)
and $\psi\geq 1$ satisfies $\frac{1}{\phi}+\frac{1}{\psi}=1$. Note that the inner product in (22) is applied to each component of $\mathcal{Q}$ and $\mathcal{V}$, where $\mathcal{V}=\bm{\mathrm{L}}\bm{\mathrm{x}}$. With these, we rewrite (21) as
$\min_{\bm{\mathrm{x}}\in\mathcal{C}}\max_{\mathcal{Q}\in B_{\infty,\psi}}\frac{1}{2}\|\bm{\mathrm{x}}-\bm{\mathrm{s}}^{k}\|_{\bm{\mathrm{P}}}^{2}+\bar{\lambda}\left\langle\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q},\bm{\mathrm{x}}\right\rangle,$ (24)
where $\bm{\mathrm{L}}^{\mathcal{T}}$ denotes the adjoint of $\bm{\mathrm{L}}$. After some rewriting, (24) becomes
$\begin{array}[]{rl}\max\limits_{\mathcal{Q}\in B_{\infty,\psi}}\min\limits_{\bm{\mathrm{x}}\in\mathcal{C}}&\|\bm{\mathrm{x}}-(\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q})\|_{\bm{\mathrm{P}}}^{2}\\ &\quad-\|\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q}\|_{\bm{\mathrm{P}}}^{2}.\end{array}$ (25)
Because $\bm{\mathrm{x}}$ appears only in the first term of (25), the minimizer of $\bm{\mathrm{x}}$ is exactly the WPM associated with $\delta_{\mathcal{C}}$:
$\bar{\bm{\mathrm{x}}}^{*}=\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q}).$ (26)
Substituting (26) into (25), we get
$\begin{array}[]{rl}\mathcal{Q}^{*}=\arg\min\limits_{\mathcal{Q}\in B_{\infty,\psi}}&\|\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q}\|_{\bm{\mathrm{P}}}^{2}\\ &\,\,-\Big{\|}\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q})\\ &\quad\quad\quad-(\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q})\Big{\|}_{\bm{\mathrm{P}}}^{2}.\end{array}$ (27)
Following [53], the gradient and the corresponding Lipschitz constant of the objective function in (27) are
$-2\bar{\lambda}\bm{\mathrm{L}}\,\mathrm{prox}_{\delta_{\mathcal{C}}}^{\bm{\mathrm{P}}}(\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q}),$ (28)
and $2\sigma^{\bm{\mathrm{P}}}_{\mathrm{min}}\bar{\lambda}^{2}\|\bm{\mathrm{L}}\|^{2}$, respectively. $\sigma^{\bm{\mathrm{P}}}_{\mathrm{min}}$ is the smallest eigenvalue of $\bm{\mathrm{P}}$. Since the objective function in (27) is differentiable, we apply APG to solve (27). After getting $\mathcal{Q}^{*}$, we compute
$\bm{\mathrm{x}}^{k+1}=\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\bm{\mathrm{s}}^{k}-\bar{\lambda}\bm{\mathrm{P}}^{-1}\bm{\mathrm{L}}^{\mathcal{T}}\mathcal{Q}^{*}).$ (29)
Both (28) and (29) require computing $\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\cdot)$ and $\bm{\mathrm{P}}^{-1}$, which can be computationally expensive for a general $\bm{\mathrm{P}}$. A natural way to choose $\bm{\mathrm{P}}$ is to approximate the Hessian $\bm{\mathrm{A}}^{\mathrm{T}}\bm{\mathrm{A}}$. By using RNP, $\bm{\mathrm{P}}$ (see (11)) can be rewritten as $\bm{\mathrm{I}}+\bar{\bm{\mathrm{U}}}\bar{\bm{\mathrm{U}}}^{\mathrm{T}}$ with $\bar{\bm{\mathrm{U}}}=\bm{\mathrm{U}}\sqrt{\frac{1}{\hat{s}_{K}+\mu}(\hat{\bm{\mathrm{S}}}+\mu\bm{\mathrm{I}})-\bm{\mathrm{I}}}$, where the square root here is applied to the diagonal elements of the inner item. Clearly, $\bm{\mathrm{P}}$ has the same structure as $\bm{\mathrm{W}}$ proposed in Theorem 1, so getting $\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\cdot)$ reduces to addressing a nonlinear equation (8). Thus, the total computation for $\mathrm{prox}^{\bm{\mathrm{P}}}_{\delta_{\mathcal{C}}}(\cdot)$ can be dramatically reduced, as demonstrated by our numerical experiments in Section IV-B. Note that ${\hat{s}}_{K}$ is the $K$th eigenvalue of $\bm{\Phi}\left<\bm{\Omega}\right>$, so the last column in $\bar{\bm{\mathrm{U}}}$ will be a zero vector. In practice, we found that using $\sqrt{{\hat{s}}_{K}}$ instead of ${\hat{s}}_{K}$ resulted in faster convergence. So, in this paper, we use ${\hat{s}}_{K}\leftarrow\sqrt{{\hat{s}}_{K}}$ in (11). Since $\bm{\mathrm{U}}$ is an orthogonal matrix, we have $\sigma^{\bm{\mathrm{P}}}_{\mathrm{min}}=1$, so the Lipschitz constant of the cost function in (27) becomes $2\bar{\lambda}^{2}\|\bm{\mathrm{L}}\|^{2}$.
When $\bm{\mathrm{L}}$ is invertible and $\mathcal{C}=\mathbb{R}^{N}$, we can rewrite (4) as
$\min_{\bar{\bm{\mathrm{x}}}\in\mathbb{R}^{\bar{N}}}\frac{1}{2}\|\bm{\mathrm{A}}{\bm{\mathrm{L}}}^{-1}\bar{\bm{\mathrm{x}}}-\bm{\mathrm{y}}\|_{2}^{2}+\lambda\|\bar{\bm{\mathrm{x}}}\|_{1,\phi},$ (30)
where $\bm{\mathrm{x}}=\bm{\mathrm{L}}^{-1}\bar{\bm{\mathrm{x}}}$. If $\phi=1$ and $\bm{\mathrm{L}}$ refers to a wavelet transform, (30) represents wavelet-regularized reconstruction. In such cases, we do not need to compute WPM iteratively, saving further computation. By setting $\bm{\mathrm{P}}$ to approximate $(\bm{\mathrm{L}}^{-1})^{\mathcal{T}}\bm{\mathrm{A}}^{\mathrm{T}}\bm{\mathrm{A}}\bm{\mathrm{L}}^{-1}$ with RNP, we only need to solve a small nonlinear equation to compute the associated WPM. We did not observe any degradation in our subsequent CT reconstruction when using (30), even though it does not account for the constraint $\mathcal{C}$.
Algorithm 3 Weighted accelerated proximal gradient (WAPG) method.
0:  Initialization $\bm{\mathrm{u}}^{1},~{}\bm{\mathrm{x}}^{1}$, stepsize $\alpha$, $t^{0}=1$, $\bm{\mathrm{P}}$, and maximum number of iterations Max_Iter.
0:  $\bm{\mathrm{x}}^{\mathrm{Max\_Iter}}$.
1:  for $k=1,2,\cdots,$ Max_Iter do
2:     $\bm{\mathrm{x}}^{k+1}=\mathrm{prox}_{(\alpha\lambda)g+\delta_{\mathcal{C}}}^{\bm{\mathrm{P}}}(\bm{\mathrm{u}}^{k}-\alpha\bm{\mathrm{P}}^{-1}\nabla f(\bm{\mathrm{u}}^{k}))$
3:     $t^{k}=\frac{1+\sqrt{1+4(t^{k-1})^{2}}}{2}$
4:     $\bm{\mathrm{u}}^{k+1}=\bm{\mathrm{x}}^{k+1}+\frac{t^{k-1}-1}{t^{k}}(\bm{\mathrm{x}}^{k+1}-\bm{\mathrm{x}}^{k})$
5:  end for
III-C On-the-Fly Nyström Preconditioner Implementation
Sections III-A and III-B show how to adapt RNP to accelerate IRM and APG for (3) and (4), respectively. However, if computing RNP is expensive, then using RNP may lose its practical utility. Indeed, in (19), $\bm{\mathrm{W}}_{f}^{k}$ and $\bm{\mathrm{W}}_{g}^{k}$ change at each iteration, so the preconditioner should be recomputed. Therefore, the entire algorithm may be slower in terms of wall time than one without using RNP if obtaining RNP is computationally expensive.
Algorithm 1 shows that, to compute RNP, one must evaluate $\bm{\Phi}\bm{\Omega}$. Each column of $\bm{\Omega}$ represents an image. In image reconstruction $\bm{\mathrm{A}}$ and $\bm{\mathrm{L}}$ are represented as operators, so the associated $\bm{\Phi}$ is also an operator, such as $\bm{\Phi}=\bm{\mathrm{A}}^{\mathrm{T}}\bm{\mathrm{A}}$ in (4) and $\bm{\Phi}=(\bm{\mathrm{L}}^{-1})^{\mathcal{T}}\bm{\mathrm{A}}^{\mathrm{T}}\bm{\mathrm{A}}\bm{\mathrm{L}}^{-1}$ in (30). For some $\bm{\mathrm{A}}$ and $\bm{\mathrm{L}}$, evaluating $\bm{\mathrm{A}}\bm{\mathrm{x}}$ and $\bm{\mathrm{L}}\bm{\mathrm{x}}$ many times could be computationally expensive. Moreover, in practice, one computes $\bm{\Phi}\bm{\Omega}$ by applying $\bm{\Phi}$ to each column of $\bm{\Omega}$ sequentially, which could be slow when $K$ is moderately large. So step 2 in Algorithm 1 will dominate the overall computation, making it crucial to significantly reduce this computational cost.
On classical CPU computational platforms, a parallel scheme may be used to evaluate $\bm{\Phi}\bm{\Omega}$. However, a central server still must collect the results to formulate $\bm{\mathrm{Y}}$, making communication time a potential bottleneck. Modern machine learning libraries (e.g., PyTorch [62] and TensorFlow [63]) support a batch mode, allowing efficient parallel computation of $\bm{\Phi}\bm{\Omega}$. Additionally, these libraries are highly optimized for modern GPU computational platforms, and the batch mode is easy to implement with basic Python knowledge. These features enable us to compute RNP on-the-fly. Table IV presents the wall time of computing RNP for CT reconstruction with different acquisitions on CPU and GPU computational platforms, clearly showing the advantage of using GPU.
TABLE I: The associated minimization problems and the used priors for different imaging tasks. $\nabla$ and $\nabla^{2}$ refer to the first- and second-difference operators, respectively.
Image Task Problem $q/\phi$ $\bm{\mathrm{L}}$
Debluring/Super-Resolution (3) with $p\in(0,1]$ $q=1$nabla, TV
Computed Tomography (30) $\phi=1$ wavelet
(4) $\phi=1$nabla, TV
(4) $\phi=1,2,\infty$ $\nabla^{2}$, HS
IV Numerical Experiments
This section studies the performance of using RNP on three different image reconstruction tasks: image deblurring and super-resolution with impulsive noise, and CT. Following [16], we use a TV regularizer for image deblurring and super-resolution tasks. For CT reconstruction, we consider wavelet, TV, and HS norm regularizers. Table I presents the associated minimization problems and regularizers for different image tasks in this part. All experiments were implemented with PyTorch and run on an NVIDIA GeForce RTX 3090.
IV-A Image Deblurring and Super-Resolution
For the image deblurring task, we first degraded the test images by convolving them with two different $9\times 9$ kernels: a uniform blur and a Gaussian blur with a standard deviation of $1.6$. In the image super-resolution task, we generated the low-resolution image by convolving the high-resolution with a $7\times 7$ Gaussian blur with a standard deviation of $1.6$, followed by downsampling with a factor of $2$. We added $5\%$ salt-and-pepper noise to the degraded images. [^2] We first randomly chose $5$% of the pixels and set them to $1$. Then, we randomly chose another $5$% of the pixels and set them to 00. The supplementary material also reports the performance of using RNP with different noise levels. We set the sketch size $K=100$ and tolerance $\epsilon=10^{-4}$ in Algorithm 1 and PCG, respectively. Similar to [12], the restoration of an RGB image was conducted by converting it to the YCbCr color-space first. Next, we applied the reconstruction algorithm to the luminance channel only and then transformed the result back to the RGB domain. For the deblurring task, we simply applied a median filter to the chroma channels for denoising. For the super-resolution task, we first used a median filter to denoise the chroma channels and then applied bicubic interpolation to upscale them to the desired resolution. Moreover, the peak signal-to-noise ratio (PSNR) values of RGB images were evaluated on the luminance channel only. We used $\ell_{p}-\ell_{1}$ to denote the method with different $p$. For the one using RNP, we added “P-”, i.e., $\bm{\mathrm{P}}-\ell_{p}-\ell_{1}$.
[…]
Figure 1: The reconstructed images for different $p$ on the image deblurring and super-resolution tasks. The PSNR value is labeled at the right bottom corner of each image. The first and fifth columns are the noisy measurement and the ground truth, respectively. The first (respectively, second) row is the reconstructed starfish (respectively, leaves) image at $10$th (respectively, $6$th) iteration for the image deblurring task with uniform (respectively, Gaussian) blur. The third row is the reconstructed bike image at $7$th iteration for the image super-resolution task with a downsampling factor $2$ and $5$% salt-and-pepper noise.
TABLE II: Comparison of PSNR and ST for different $p$ in the image deblurring task with uniform and Gaussian blurs. We ran IRM $20$ iterations on each image and presented the highest PSNR values within these $20$ iterations along with the associated ST values. The highest PSNR and ST values for each image are marked in bold.
Image Butterfly Boats C. Man House Parrot Barbara Starfish Peppers Leaves
Deblurring: Uniform blur, 5% salt-and-pepper noise
PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST
$p=1$ $31.6$ $0.94$ $32.9$ $0.94$ $31.8$ $0.95$ $35.9$ $0.93$ $32.3$ $0.95$ $27.9$ $0.95$ $30.8$ $0.94$ $32.2$ $0.94$ $30.7$ $0.94$
$p=0.8$ $35.3$ $0.96$ $35.9$ $0.95$ $35.1$ $0.97$ $38.7$ $0.96$ $35.7$ $0.97$ $31.1$ $0.97$ $34.2$ $0.96$ $35.2$ $0.96$ $34.7$ $0.96$
$p=0.5$ $\bm{39.9}$ $\bm{0.98}$ $\bm{39.9}$ $\bm{0.97}$ $\bm{38.8}$ $\bm{0.98}$ $\bm{42.5}$ $\bm{0.98}$ $\bm{39.2}$ $\bm{0.98}$ $\bm{34.9}$ $\bm{0.98}$ $\bm{38.4}$ $\bm{0.98}$ $\bm{38.8}$ $\bm{0.98}$ $\bm{39.7}$ $\bm{0.98}$
Deblurring: Gaussian blur, 5% salt-and-pepper noise
$p=1$ $29.5$ $0.95$ $30.9$ $0.96$ $29.4$ $0.97$ $34.1$ $0.96$ $31.9$ $0.96$ $28.1$ $0.97$ $29.9$ $0.96$ $31.3$ $0.95$ $29.7$ $0.96$
$p=0.8$ $31.5$ $0.98$ $32.9$ $0.98$ $31.5$ $0.98$ $35.6$ $0.98$ $33.5$ $0.98$ $30.3$ $0.98$ $31.9$ $0.98$ $32.5$ $0.98$ $31.5$ $0.98$
$p=0.5$ $\bm{33.5}$ $\bm{0.99}$ $\bm{35.2}$ $\bm{0.99}$ $\bm{33.5}$ $\bm{0.99}$ $\bm{37.4}$ $\bm{0.99}$ $\bm{34.9}$ $\bm{0.99}$ $\bm{31.9}$ $\bm{0.99}$ $\bm{33.7}$ $\bm{0.99}$ $\bm{34.1}$ $\bm{0.99}$ $\bm{33.3}$ $\bm{0.99}$
TABLE III: Comparison of PSNR and ST for different $p$ in the image super-resolution task. We ran IRM $20$ iterations on each image and presented the highest PSNR values within these $20$ iterations along with the associated ST values. The highest PSNR and ST values for each image are marked in bold.
Super-resolution: scaling = 2, 5% salt-and-pepper noise
Image Butterfly Flower Girl Parth. Parrot Racoon Bike Hat Plants
PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST PSNR ST
$p=1$ $25.6$ $0.74$ $27.4$ $0.74$ $31.6$ $0.78$ $24.5$ $0.43$ $27.5$ $0.76$ $27.3$ $0.68$ $23.2$ $0.76$ $29.4$ $0.76$ $31.9$ $0.76$
$p=0.8$ $27.1$ $0.94$ $28.7$ $0.93$ $32.5$ $0.91$ $25.4$ $0.91$ $28.6$ $0.94$ $28.4$ $0.90$ $24.3$ $0.93$ $30.3$ $0.94$ $33.3$ $0.92$
$p=0.5$ $\bm{28.1}$ $\bm{0.96}$ $\bm{29.6}$ $\bm{0.97}$ $\bm{33.1}$ $\bm{0.96}$ $\bm{25.9}$ $\bm{0.96}$ $\bm{29.7}$ $\bm{0.97}$ $\bm{29.4}$ $\bm{0.97}$ $\bm{25.3}$ $\bm{0.97}$ $\bm{31.4}$ $\bm{0.96}$ $\bm{34.4}$ $\bm{0.96}$
IV-A1 Image Deblurring
(a) Uniform blur.
(b) Uniform blur.
(c) Gaussian blur.
(d) Gaussian blur.
Figure 2: PSNR values versus iteration and wall time for different $p$ in the image deblurring task. First (respectively, second) row is tested on the starfish (respectively, leaves) image.
(a) Starfish & uniform blur.
(b) Leaves & Gaussian blur.
Figure 3: Number of iterations within CG for different $p$ in the image deblurring task. The legend is identical to Figure 2.
Figure 2 presents the PSNR values versus iteration and wall time for different $p$. The first (respectively, second) row of Figure 2 was tested on the starfish (respectively, leaves) image with uniform (respectively, Gaussian) blur. Figures 2(a) and 2(c) show a small $p$ yielded a higher PSNR than a large $p$. It is not supervised because a small $p$ is more robust to outliers [64]. However, (3) becomes nonconvex for $p<1$, so solving (3) is more challenging. Indeed, Figures 2(b) and 2(d) illustrate that a small $p$ required much more wall time than a large one. Figure 2 illustrates that using RNP significantly accelerated the convergence speed in terms of wall time. The time for computing RNP was included in the whole wall time. From Figures 2(b) and 2(d), we even saw $\mathrm{P}-\ell_{0.5}-\ell_{1}$ converged faster than $\ell_{1}-\ell_{1}$ in terms of the wall time illustrating the effectiveness of using RNP. Figure 3 describes the number of iterations within CG at each iteration of IRM for different $p$, with the uniform and Gaussian kernels for the starfish and leaves images, respectively. Using RNP reduced the number of iterations required by CG more than $90$%, while achieving the same solution accuracy. The first and second rows of Figure 1 present the noisy, reconstructed, and ground truth images, where the degradation is reduced significantly.
To quantitatively measure the performance of using RNP, we define a “saved time” (ST) criterion:
$\mathrm{ST}=\frac{\mathrm{Time}_{\text{w/o}}-\mathrm{Time}_{\text{w}}}{\mathrm{Time}_{\text{w/o}}},$ (31)
where $\mathrm{Time}_{\text{w}}$ and $\mathrm{Time}_{\text{w/o}}$ denote the wall time with and without using RNP, respectively. If $\mathrm{ST}$ is close to $1$, then using RNP significantly reduced the wall time. To study the performance of RNP further, we tested on additional $8$ images. We ran IRM $20$ iterations for different $p$. Table II presents the highest PNSR and the associated ST value for each image within $20$ iterations. Using RNP saved more than $95\%$ wall time than without RNP, illustrating the effectiveness and efficiency of using RNP. Moreover, Table II also shows that a smaller $p$ yielded a higher PSNR.
IV-A2 Image Super-Resolution
(a) Bike.
(b) Bike.
Figure 4: PSNR values versus iteration and wall time for different $p$ in the image super-resolution task.
Figure 5: Number of iterations within CG for different $p$ in the super-resolution task. The downsampling factor is $2$. The legend is identical to Figure 4.
Figure 4 shows the PSNR values for different $p$ in the image super-resolution of the bike image. The results are consistent with the trends in Section IV-A1, where a small $p$ yielded a higher PSNR than a larger one and using RNP significantly accelerated the convergence speed in terms of wall time. Moreover, Figure 5 presents the number of CG iterations for different $p$; the preconditioned one required fewer iterations illustrating the effectiveness of using RNP. The third row of Figure 1 presents the reconstructed images, where we observed that a small $p$ yielded a higher quality image. We tested on additional $8$ images to study the performance of using RNP further. Table III shows using RNP saved almost $70\%$ (respectively, $95\%$) time for $p=1$ (respectively, $p=0.5$) in line with the observation in Section IV-A1.
IV-A3 The Choice of Sketch Size
Figure 6: Comparison of different $K$ in the image deblurring task with Gaussian kernel on the leaves image. (a): the number of iterations within CG in each iteration; (b): PSNR values versus wall time.
From Theorem 2, one option for choosing the sketch size is to set $K=2\lceil 1.5d_{\mathrm{eff}}(\mu)+1\rceil$. However, $d_{\mathrm{eff}}(\mu)$ is difficult to compute in practice. Moreover, $d_{\mathrm{eff}}(\mu)$ can be extremely large, making computing RNP time- and memory-intensive, even with an on-the-fly implementation. Indeed, for an extremely large $K$, steps 5-7 in Algorithm 1 will dominate the computation instead of step 2. Furthermore, applying RNP can be also expensive for an extremely large $K$. Murray et al. [65] presented an adaptive strategy to update $K$ at each iteration, avoiding the need for $d_{\mathrm{eff}}(\mu)$. However, the adaptive strategy requires calling Algorithm 1 multiple times and executing many iterations of the power method to estimate the error between $\bm{\Phi}$ and $\bm{\mathrm{U}}\hat{\bm{\mathrm{S}}}\bm{\mathrm{U}}^{\mathrm{T}}$. Therefore, it is unsuitable for our experimental settings, where RNP must be computed on-the-fly. We found that using a fixed, moderate $K$ works well in practice.
Figure 6 presents the number of iterations within CG in the image deblurring task with Gaussian blur on the leaves image for different $K$. For $K>20$, the number of iterations within CG did not change significantly. In general, a larger $K$ would yield a more effective preconditioner, but it also increases the computational cost to compute and apply. Figure 6 shows the reconstructed PSNR values versus wall time, where clearly $K=300$ required more wall time than the others. Moreover, $K=50$ and $K=100$ performed better than the others. Since $K=100$ performed slightly better than $K=50$, we simply set $K=100$ in this subsection.
IV-B Computed Tomography Reconstruction
We studied the performance of using RNP for CT reconstruction with three different regularizers: wavelet, TV with $\phi=1$, and HSϕ norm. For HSϕ, we mainly studied $\phi=1$. We used the “daub4” wavelet with $4$ levels in the “pywt” [66] toolbox. The Operator Discretization Library (ODL) [67] was used for the CT forward model. We investigated parallel-beam and fan-beam acquisition geometries with $100$ projection views. We scaled two slices from subject “067” in the AAPM CT Grand Challenge data [68] to $[0,1]$ as the test images; see Figure 7. Due to the page limit, the supplement shows the reconstruction results for Figure 7 and for the fan-beam geometry. Again we use “w/o” to denote results without using RNP, while “w-$K$” denotes RNP with sketch size $K$. Table IV summaries the wall time for computing RNP for different acquisitions with various regularizers on both CPU and GPU computational platforms, clearly highlighting the advantage of using the batch mode and the GPU computational platforms. In general, a larger $K$ would yield a faster convergence in terms of iterations than a small $K$. However, large $K$ would also increase the computation of applying RNP. Here we set $K=20$ for wavelet and TV-based reconstruction, and $K=100$ for HS1-based reconstruction, because these values worked well in practice.
Figure 7: Ground truth CT images.
TABLE IV: Comparison of wall time for obtaining RNP in computed tomography reconstruction on CPU and GPU.
Parallel beam Fan beam
Wavelet TV HS Wavelet TV HS
$K$ $20$ $20$ $100$ $20$ $20$ $100$
GPU $0.11$s $0.10$s $0.40$s $0.11$s $0.09$s $0.46$s
CPU $0.48$s $0.40$s $4.20$s $0.48$s $0.52$s $4.25$s
Figure 9 presents the comparison of using RNP for wavelet based CT reconstruction. Figures 9 and 9 shows that w-$20$ converged faster than w/o in terms of iterations, illustrating the effectiveness of using RNP. Moreover, Figures 9 and 9 shows that w-$20$ converged faster than w/o in terms of wall time, demonstrating the wall time for solving the related WPM with Theorem 1 is negligible. Figure 10 presents the results of TV-based reconstruction that shows similar trends as in wavelet-based regularization. Figure 8 illustrates the reconstructed images at iterations $10$, $20$, $40$, and $60$, for both w-$K$ and w/o.
Refer to caption10101010w/o23.623.623.623.6Refer to caption27.227.227.227.2wRefer to caption20202020w/o28.228.228.228.2Refer to caption32.332.332.332.3wRefer to caption40404040w/o33.233.233.233.2Refer to captionw34.334.334.334.3Refer to caption60606060w/o34.234.234.234.2Refer to captionw34.734.734.734.7WaveletRefer to caption23.623.623.623.6Refer to caption25.925.925.925.9Refer to caption28.428.428.428.4Refer to caption31.231.231.231.2Refer to caption33.233.233.233.2Refer to caption33.633.633.633.6Refer to caption33.733.733.733.7Refer to caption33.733.733.733.7TVRefer to caption23.623.623.623.6Refer to caption29.729.729.729.7Refer to caption28.228.228.228.2Refer to caption33.933.933.933.9Refer to caption33.433.433.433.4Refer to caption36.236.236.236.2Refer to caption35.135.135.135.1Refer to caption37.037.037.037.0HS1
Figure 8: The parallel-beam CT reconstructed images with wavelet, TV, and HS1 based regularization at iterations $10$, $20$, $40$, and $60$. Columns $1$, $3$, $5$, and $7$ (respectively, $2$, $4$, $6$, and $8$) show the reconstructions without (respectively, with) RNP. The associated PSNR values are listed at the right corner of each image.
Figure 9: Comparison of using RNP for wavelet based CT reconstruction. w/o denotes the one without using RNP.
Figure 10: Comparison of using RNP for TV based CT reconstruction. w/o denotes the one without using RNP.
Figure 11: Comparison of using RNP for HS1 based CT reconstruction. w/o denotes the one without using RNP.
Since TV based reconstruction will introduce blocking artifacts [33], we included the study of HS norm based reconstruction. Figure 11 shows the results of HS1 based reconstruction, where we observed that using RNP accelerated the convergence compared to not using RNP. Moreover, we observed that HS1 yielded a higher PSNR than wavelet and TV regularizers. The supplementary material reports the performance of HS2 and HS based reconstruction that trended similar to HS1.
V Conclusion and Future Work
In this paper, we introduced RNP to accelerate variational image reconstruction. Additionally, we showed how to efficiently adapt RNP to solve reconstruction problems that involve nonsmooth priors. By leveraging modern GPU computational platforms and PyTorch, we achieved an on-the-fly implementation, making RNP suitable for applications that require real-time processing. We extensively evaluated the performance of RNP on image deblurring, super-resolution tasks, and computed tomography with various regularizers, demonstrating its effectiveness and efficiency in accelerating variational image reconstruction.
Extending RNP to magnetic resonance imaging (MRI) reconstruction presents an interesting future direction. Modern MRI scanners use multi-coils to acquire k-space data, which causes the forward model to change for each scan, requiring the preconditioner to be computed on-the-fly. Additionally, in functional MRI applications, the blood oxygenation level-dependent (BOLD) contrast, which is sensitive to magnetic field inhomogeneity, is primarily used. In some cases, images are jointly reconstructed with the associated field maps [69], causing the forward model to change at each iteration. Given RNP’s on-the-fly implementation, we believe it holds great potential as a preconditioner for acceleration.
Appendix A Examples of $\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{1,\phi}$
Let $\bm{\mathrm{X}}\in\mathbb{R}^{N_{1}\times N_{2}}$ be the matrix form of $\bm{\mathrm{x}}$, where $\bm{\mathrm{x}}$ is the column stacking of $\bm{\mathrm{X}}$. So if $\bm{\mathrm{L}}$ represents the first-order differential operator, we have $G=N$ and
$\bm{\mathrm{v}}_{l}=\begin{bmatrix}\bm{\mathrm{X}}_{i,j}-\bm{\mathrm{X}}_{i-1,j}\\ \bm{\mathrm{X}}_{i,j}-\bm{\mathrm{X}}_{i,j-1}\end{bmatrix},$
with $l=j\,N_{1}+i$ and $\bm{\mathrm{X}}_{i,j}$ denotes the component of $\bm{\mathrm{X}}$ at the $i$th row and the $j$th column. So for $\phi=1$ (respectively, $\phi=2$), we have $\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{1,1}=\sum_{l}|\bm{\mathrm{X}}_{i,j}-\bm{\mathrm{X}}_{i-1,j}|+|\bm{\mathrm{X}}_{i,j}-\bm{\mathrm{X}}_{i,j-1}|$ (respectively, $\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{1,2}=\sum_{l}\sqrt{(\|\bm{\mathrm{X}}_{i,j}-\bm{\mathrm{X}}_{i-1,j})^{2}+(\bm{\mathrm{X}}_{i,j}-\bm{\mathrm{X}}_{i,j-1})^{2}}$), which represents the anisotropic (respectively, isotropic) total variation.
If $\bm{\mathrm{L}}$ represents the second-order finite-difference operator, we have
$\bm{\mathrm{v}}_{l}=\begin{bmatrix}v_{l}^{11}&v_{l}^{12}\\[5.0pt] v_{l}^{21}&v_{l}^{22}\end{bmatrix},$
with $l=j\,N_{1}+i$ and $v_{l}^{11}=\bm{\mathrm{X}}_{i-1,j}-2\bm{\mathrm{X}}_{i,j}+\bm{\mathrm{X}}_{i+1,j},\,v_{l}^{22}=\bm{\mathrm{X}}_{i,j-1}-2\bm{\mathrm{X}}_{i,j}+\bm{\mathrm{X}}_{i,j+1},\,v_{l}^{12}=v_{l}^{21}=\frac{1}{4}\big{(}\bm{\mathrm{X}}_{i+1,j+1}-\bm{\mathrm{X}}_{i+1,j-1}-\bm{\mathrm{X}}_{i-1,j+1}+\bm{\mathrm{X}}_{i-1,j-1}\big{)}$. Thus $\|\bm{\mathrm{L}}\bm{\mathrm{x}}\|_{1,\phi}=\sum_{l}\|\bm{\mathrm{v}}_{l}\|_{\phi}$.